OpenStudy (anonymous):

I need someone to check my work! What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = –51 ? ANSWER: -123

OpenStudy (cwrw238):

35th term = a1 + 34d

OpenStudy (cwrw238):

find d by :- a17 - a1 / 16 = -51 - 13 / 16 = -4

OpenStudy (anonymous):

okay.

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

that gives 13 - 136 = -123

OpenStudy (anonymous):

ty.

OpenStudy (cwrw238):

no pobs